OpenStudy (anonymous):

Write the equation of the line, in standard form, that has a y-intercept of 2 and is parallel to 2x + y = -5.

OpenStudy (anonymous):

@eta

OpenStudy (anonymous):

See, lines are parallel, again find slope of given line..

OpenStudy (anonymous):

Slope is -2

OpenStudy (anonymous):

So, the other line will have same slope and its y-intercept is also given directly here, so what is the line's equation?

OpenStudy (anonymous):

\[y = (Slope) x + (y-intercept)\]

OpenStudy (anonymous):

y=-2x+2

OpenStudy (anonymous):

Reduced my work.. :P

OpenStudy (anonymous):

Good.. That is it..!!

OpenStudy (anonymous):

But that's not standard form though

OpenStudy (anonymous):

@eta

OpenStudy (anonymous):

Can't you write that in standard form?

OpenStudy (anonymous):

See, what is standard form? It is shown below: \[ax + by = c\]

OpenStudy (anonymous):

x and y terms on left hand side constant term on right hand side The coefficient of \(x\) ie \(a\), should be positive..

OpenStudy (anonymous):

Add \(2x\) both the sides and you will have your standard form dear..!!

OpenStudy (anonymous):

2x+y=2 ..?

OpenStudy (anonymous):

Or is it 2x-y=2?

OpenStudy (anonymous):

Nope, look it once again..

OpenStudy (anonymous):

How you got \(-y\) ?? How you ended up with \(-\) sign with \(y\) ?

OpenStudy (anonymous):

Is it 2x+y=2?

OpenStudy (anonymous):

Good.. :)